AutoPartsWarehouse.com purchases chrome rims for $720 (for a set of 4) and uses a markup of 65% on selling price. What will be the percent markup on cost? Report your answer as a a percentage and round to the nearest percent.
Answer #1

Ans 186%

Let the selling price be x
Cost + Profit = Selling Price
720 + 65%*x = x
720 = x - 0.65x
720 = 0.35 x
x = 720 / 0.35
x = 2057.142857
Profit = 65% * x
65% * 2057.142857
1337.1428571
Markup on Cost = Profit / Cost * 100
1337.1428571 / 720 * 100
185.71%